### v1

# Investigating Experiential Effects in Online Chess using a Hierarchical Bayesian Analysis

Abstract
The presence or absence of winner-loser effects is a widely discussed phenomenon across both sports and psychology research. Investigation of such effects is often hampered by the limited availability of data. Online chess has exploded in popularity in recent years and provides vast amounts of data which can be used to explore this question. With a hierarchical Bayesian regression model, we carefully investigate the presence of such experiential effects in online chess. Using a large quantity of online chess data, we see little evidence for experiential effects that are consistent across all players, with some individual players showing some evidence for such effects. Given the challenging temporal nature of this data, we discuss several methods for assessing the suitability of our model and carefully check its validity.
1 Introduction
It is rare that the success or failure of an individual competitor or sports team is not discussed in the context of their recent performance. The idea of momentum in competition, broadly defined, has been examined repeatedly in the scientific literature. In sports analytics this is often investigated in the context of the ``hot hand'' phenomenon, where players experience streaks of success that appear unlikely to occur due to chance [34, 38, 40]. Similarly, other studies have directly considered the question of momentum in sport, both within and between games [20, 46].
A large body of research in the psychology and animal behavior literature explores a specific formulation of this concept of momentum. It is commonly referred to as winner-loser or experiential effects [^1] We will use these two names interchangeably throughout this work., defined as success (or failure) in recent competition leading to an increased probability of success (or failure) in future competition [12]. The presence of these winner-loser effects have been hypothesized to facilitate the formation of stable dominance hierarchies [12, 27]. In animals, loser effects appear to consistently be stronger than winner effects (e.g. sticklebacks [3]; crayfish [6]; copperhead snakes [42]). In humans, several studies have revealed some evidence for winner and loser effects in competitive sports including tennis, football and judo [8, 20, 21, 39]. For example, winners of tennis tie-breaks are more likely to win subsequent sets when players are evenly matched [39]. Winner and loser effects have also been experimentally demonstrated in humans playing competitive video games [44]. However, there is still considerable discussion about the existence and consistency of these effects across individuals, contexts and species.
Indeed, there are still several outstanding key questions regarding the presence of such experiential effects, including
  • Are winner/loser effects strongest when considering the immediate previous outcome or several previous outcomes?
  • Do the size of winner/loser effects vary across rating levels such as between players/teams of stronger versus lesser ability?
  • Are winner/loser effects stable across time?
  • Do winner/loser effects vary within individuals across contexts such as across different formats of competitions?
One of the main challenges posed by these questions is the availability of appropriate data to investigate these phenomena. Online chess has enjoyed widespread popularity in recent years, with millions of regularly active users on several major sites. The most common format of chess played on these sites is ``Standard'' rated games. These games are often short ``bullet'' games, lasting less than three minutes. It is common for players to play many such games in succession. In these games, players are randomly paired with another user currently wishing to play at the same time and of similar estimated ability. The color that each player is assigned (White or Black) is randomly chosen. Games of this form provide an excellent setting to investigate the presence of psychological momentum and experiential effects in such sports and can be used to consider the key questions posed above. Any individual plays a wide range of players, who may utilise different styles and strategies of chess. The random assignment of opponents and individual game settings, coupled with the vast amount of games many users play, makes online chess an excellent setting to investigate the presence of winner-loser effects in a controlled environment.
In this paper we use a Bayesian hierarchical logistic regression model to investigate the presence of winner-loser effects in online chess, using publicly available data. We carefully motivate and describe the model chosen. Then, we use this model to examine if and when these effects occur, and how they vary across player ability and game format. While we do not believe this model is a complete representation of the process underlying these games, we use multiple model checking procedures to ensure our model is reasonable.
We find little evidence for strong persistent winner-loser effects, with this conclusion consistent across the range of players and abilities analysed. Individuals appear to show some variability, and there is some evidence that some players demonstrate such effects. However, these effects do not appear to be systematic.
This article is organized as follows. We first provide a brief history of the study of these effects across disciplines in Section 1.1. In Section 2 we describe in detail the data we use here, highlighting key properties which make it well suited for the study of winner-loser effects. Section 3 gives the proposed hierarchical Bayesian regression model, detailing the model and prior specification. In Section 4 we fit the proposed model, highlighting the results for two cohorts of chess players of differing abilities. Section 5 provides a detailed discussion of several procedures used to ensure our model is able to correctly capture the temporal structure present in such data. Finally, in Section 6 we summarise the results of this model and highlight important considerations for future work.
1.1 A history of the study of experiential effects in human and animal competition
Winner-loser effects have consistently been demonstrated across species engaged in various competitive interactions [48, 43, 28, 35, 25, 31, 36, 1, 26, 37, 4, 13, 47, 30]. The most common experimental design utilizes a series of dyadic matchups whereby individuals are induced to win or lose against weaker or stronger opponents and are then evaluated in subsequent competitions. These studies have demonstrated that loser effects are typically more long lasting than winner effects [48, 28, 1, 47, 30]. The likelihood of winning subsequent encounters has also been shown to increase following unresolved aggressive encounters (e.g. East African cichlid fish [10]) or even following recent losses (olive fruit fly [5]). Further, variation in competitive contexts or individual variability in skill level of fighting [7], growth rate [33], age [16], and life experience with fighting [29] can all influence the strength of winner-loser effects. Functionally, loser effects may serve to reduce aggression intensity within groups, with fewer attacks received by previous losers, resulting in fewer injuries to losers [47, 30]. Winner effects serve to protect dominants from excessive energy usage by resolving aggressive encounters more efficiently [30]. Several theoretical models have demonstrated that winner-loser effects are sufficient to lead to the formation of stable dominance hierarchies [12, 14].
There exists relatively less evidence for the presence of winner-loser effects in human competitive interactions despite their existence being strongly hypothesized. In humans, these effects are also referred to as psychological momentum which is defined as a change in the probability of an outcome occurring as a function of the outcome of preceding events [2]. For example, it has been reported that male judoists are more likely to win their next contest if they won their previous event [9], and male tennis players are more likely to win the next point if they won the preceding point [21]. Male tennis players are also more likely to win a subsequent set against an evenly matched opponent after winning a close tiebreak [39]. These effects were not observed in female players. Similar instances of psychological momentum have been observed from large datasets of competitive sports, including that hitting the bullseye in elite archery improves performance on the next shot [50], and that the likelihood of winning double-headers in softball and baseball increases if both games are played on the same day rather than over multiple days [15, 19]. Alternatively, researchers have experimentally manipulated the outcomes of competitive interactions in video games and shown that individuals who are randomly assigned to be winners perform better in future contests than do those who are randomly assigned to be losers [44].
Notably, there has been a relative lack of investigation of human competitive interactions that incorporate thousands of competitive interactions over a longer period of time such as is available with data from online chess matches. Similarly, the existing models for examining such effects are often limited or somewhat domain specific. In this work we develop a natural statistical model for such interaction data which can be applied to any dataset of repeated competition. As discussed below, it can also incorporate information about these competitions as available.
2 Data
In this work we use rated standard games played on Lichess, a popular open source chess server. Lichess and other chess platforms pair random opponents who are believed to have a similar chess ability, to create games where both players have an opportunity to win. In online and over the board chess, it is common to estimate the ability of an individual using a numeric rating based on their past performance. When an individual creates an account on Lichess, they are a given a default rating of 1500 across all chess formats. As the player plays rated games at a specific time format, their rating will update based on their performance and the rating of their opponents. Lichess uses the Glicko-2 rating system [24], which also estimates a measure of variability of this rating. This variability decreases as more games are played and the estimated rating becomes a more accurate measure of a players ability at that time format [23]. In Figure 1 we show the rating evolution of two players with a similar number of games played. We see that in each case their rating improves over time, with significant variability before somewhat stabilizing.
Lichess allows a wide range of variants and time formats of chess, with games ranging from several minutes to several days. The only variant we consider are standard chess formats, which is the vast majoirty of games on Lichess. In these time controlled chess games, each player starts with the same amount of time to make their moves. The two most common formats are ``Bullet'' (games where each player has 1-2 minutes) and ``Blitz'' (games where each player has 3-5 minutes). These short games commonly end with one player winning via checkmate, resignation or running out of time. Draws are possible but relatively rare in these formats. We consider only the most played time control in bullet and blitz for analysis: 60+0 (60 seconds for all moves with no increment) and 180+0 (3 minutes for all moves with no increment) respectively.
We wish to examine the presence of experiential effects across a range of chess abilities and playing history. Using a players current rating, we collected data for 40 players in 3 seperate rating cohorts each: 1700-1900, 2000-2200 and 2300-2500. Between October 23-31, 2023, we randomly selected 40 players within each of these rating ranges that had recently played bullet games of any format, and scraped information about the games they had played up to that date. These rating cohorts correspond to a range of moderate to high chess ability [^2] Lichess provides a weekly distribution of ratings on the platform. At the time of writing, 47% of bullet players had a rating lower than 1500, while only 1.5% had a rating higher than 2500.. We also collect all games played by 25 grandmaster (GM) accounts on Lichess in early 2024. Grandmaster is an official title awarded to players who demonstrate excellent ability in chess. At the time of writing, this extremely selective title has only been awarded to 2087 players in history. For the GMs we selected, the average rating was 3055 for Bullet and 2854 for Blitz games. We denote the 40 players in each rating cohort and the 25 GMs as ``focal'' players in this paper.
For every game by each focal player in the 60+0 and 180+0 time controls, we observe their current rating, the rating of their opponent, the colour of each player and the outcome of the game. For brevity, we focus our findings on the 1700-1900 and GM cohorts in what follows. The results obtained for the other cohorts, unless noted, agree with those presented below, with some of the corresponding analyses included in the supplementary materials.
2.1 Data Summary
For the 1700-1900 cohort, 36 focal players played over 590,000 60+0 Rated Bullet Games, more than 87% of all rated bullet games they play, with each playing on average over 16000 games. For the games played by grandmasters, 75% of all bullet games are 60+0, corresponding to more than 220,000 games, with each of the 25 players playing an average of almost 9000 games. For blitz games, we have 107,000 180+0 Rated Blitz Games in the 1700-1900 cohort (58% of all blitz games), with 25 focal players playing an average of 4140 games. For the GM cohort we have 33,000 games (98% of all blitz games), with each playing on average 1292 games.
Fig. 1: Evolution of Lichess Bullet ratings for one player in the 1700-1900 cohort and another in the GM cohort.
The time scale of experiential effects
A key question in the study of experiential effects is if and how these effects persist over time. In the context of teams sports such as football, there may be several days between subsequent games, which could interfere with such effects. Due to the short time format most popular in online chess, it is common for players to play many games in quick succession. For bullet and blitz games where the players do not receive an increment, the maximum game length is simply the sum of the time both players start with.
Fig. 2: Time gaps between successive bullet games (left) and the distribution of the lengths of a session of bullet games (right) for the GM cohort.
In Figure 2 we show the time between the start of successive bullet games in the GM cohort on a log scale. The maximum length of these games is 2 minutes. Across all grandmaster games, 88.6% of the games are played within 5 minutes of the previous game. This indicates that periods where players play more than one game in a row are very common (for the 1700-1900 cohort this is 85% of all bullet games).
Based on this information we define bullet games played in the same session as those which start within 5 minutes of the start of the previous game. With this definition we can then determine how many games are played in a given session. We plot the distribution of the number of games in a session, from the GM cohort, in Figure 2. The median number of games played in such a session is 5, with 78% of all sessions involving more than 1 game. It is very common for an individual to play multiple games in quick succession. It is commonly thought that experiential effects are short term, with the result of the previous competition having an impact on the next [17]. If experiential effects do occur, it seems plausible that they would be detected in games of this form, where players who win or lose can almost immediately enter another contest.
While it is believed that winner-loser effects occur based on the most recent history, we wish to investigate if longer term effects can exist also. In the model described below we will investigate the presence of these effects based on the result immediately proceeding the current game only. However, we will also investigate if these effects have a longer duration by using the performance in the previous 10 games as a predictor of the current outcome. It is still relatively common for players to play 10 or more games in succession, with 26% of sessions played by GMs, as we have defined them, containing more than 10 games [^3] We see similar proportions for the 1700-1900 cohort.. This will therefore often include periods where players have a gap between successive games. Fitting our proposed model taking account of this longer history allows investigation of more prolonged experiential effects.
3 Modeling Winner-Loser Effects with a Hierarchical Logistic Regression
We have large volumes of games where players are randomly assigned to opponents of similar ability and we observe potential confounding factors such as this difference in ability and the colour each player plays. This makes this data particularly well suited to investigate the role of experiential effects on game results.
To investigate these effects, we model the outcome of each game played by the focal players in our data using a Bayesian hierarchical logistic regression model. The model we consider is general and can readily be applied to many forms of data consisting of repeated competition over time, identifying both individual and population level effects in such data. The Bayesian nature of our model provides natural uncertainty quantification in our estimates along with tools for checking the validity of our model.
The probability a focal player wins an individual game is given by
$P(y_{ij}=1)=p_{ij}=\frac{1}{1+\exp(-(\alpha_{j}+\beta_{j}\tilde{x}_{ij}^{n}+\bm{\gamma}^{T}\bm{z_{ij}}))}.$ (1)
Here $p_{ij}$ is the probability of the $j$-th focal player winning their $i$-th game, $y_{ij}$. $\alpha_{j}$ is a player level random effect, indicating a players individual ability to win a game against an equally rated opponent when playing with the black pieces. $\beta_{j}$ is a player level random effect, accounting for the win ratio $\tilde{x}_{ij}^{n}$ of the focal player over their previous $n$ games. We discuss this in more detail below. $\bm{\gamma}$ is a vector of fixed effects corresponding to two individual game covariates, the colour played by the focal player and the difference in rating between the focal player and their opponent. We partially pool the $\alpha$ and $\beta$ coefficients with a bivariate normal prior distribution and place independent normal priors on the fixed effects. The prior specification for all parameters is given by
$\begin{pmatrix}\alpha_{j}\\ \beta_{j}\end{pmatrix}\sim\mathcal{N}(\bm{\mu},\Sigma),$ (2)
and
$\gamma_{1}\sim\mathcal{N}(0,\sigma_{g_{1}}^{2}),\ \gamma_{2}\sim\mathcal{N}(0,\sigma_{g_{2}}^{2})$
where we have
$\bm{\mu}=\begin{pmatrix}0\\ \mu_{\beta}\end{pmatrix}\mbox{ and }\Sigma=\begin{pmatrix}\tau_{1}&0\\ 0&\tau_{2}\end{pmatrix}\Omega\begin{pmatrix}\tau_{1}&0\\ 0&\tau_{2}\end{pmatrix}.$ (3)
We then place independent hyperpriors on each of these parameters, with
$\mu_{\beta}\sim\mathcal{N}(0,\sigma^{2}),$
$\tau_{1},\tau_{2},\sigma,\sigma_{g_{1}},\sigma_{g_{2}}\sim\mathcal{N}_{+}(0,1)$
and
$\Omega\sim\mbox{LKJcorr}(2).$
We normalize the win ratio by the average win ratio of player $j$ across all games they play in the dataset. If $x_{ij}^{n}$ is the proportion of games that the $j$-th player has won in the past $n$ games then $\tilde{x}_{ij}^{n}=x_{ij}^{n}-\bar{x}_{j}$, where $\bar{x}_{j}$ is their overall win proportion for games of that format. $\beta_{j}$ captures how a players win probability changes as their recent performance deviates from their overall average win ratio.
For $\bm{z_{ij}}$ we incorporate 2 covariates, namely the colour played by the focal player in an individual game and the rating difference between the focal player and their opponent. $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ where $\gamma_{1}$ describes how the win probability changes in going from playing as black to playing as white. $\gamma_{1}>0$ would indicate that against an equally rated opponent, there would be an increased probability of winning when playing as white compared to playing as black. $\gamma_{2}$ describes the relationship between win probability and the difference in rating between the focal player and their current opponent. $\gamma_{2}>0$ would indicate that the more highly rated a focal player is than their opponent, the higher the probability they would win a match, accounting for the colour assigned to each player.
We decompose the covariance matrix $\Sigma$ into a correlation matrix $\Omega$ and the individual scale of the two random effects. For the correlation matrix we choose an LKJ prior with shape parameter 2 [32]. This is a prior over the set of correlation matrices with the chosen shape parameter favoring less correlation. We use independent Inverse-Gamma priors with shape and scale 1 for all scale parameters. In initial model fitting we found that less informative priors for the scale parameters gave equivalent results while making posterior convergence more challenging.
As we are interested in identifying potential global winner-loser effects, we place a prior on $\mu_{\beta}$, the global winner-loser effect. The aim of inferring this parameter is to identify if there is evidence for a common effect across all players, rather than isolated to individual players. $\mu_{\beta}>0$ would indicate the presence of a cohort level winner-loser effect. If the focal player had won (lost) more recent games than their overall win proportion, this would lead to their probability of winning the next game increasing (decreasing). $\mu_{\beta}<0$ would indicate a reverse relationship, where poor recent performance would lead to an increased win probability in future games. We note that this model gives equal weight to both these winner and loser effects and cannot identify them separately, an important natural extension. This model can then be fit for various choices of $n$, the number of previous games considered in $\tilde{x}_{ij}^{n}$. As discussed above, we will consider $n=1$ and $n=10$, investigating both immediate and longer term experiential effects. In the appendix we demonstrate similar results with $n=5$ also.
This model is a natural approach to investigating this phenomenon in online chess, and is similar to models which have been considered in both the sport analytics and animal behaviour literature [11, 17]. The flexibility of this model allows it to be applied to a range of potential datasets consisting of repeated competition observed for a number of individuals.
4 The Fitted Model
We fit the proposed Bayesian model with Stan [45], using the CmdStanR [18] interface in R [41]. We use default settings for fitting these models, running 4 chains each with 1000 warmup and 1000 sampling iterations. We fit the proposed model for each of the four cohorts discussed previously. In each case the models indicate posterior convergence, as indicated by $\hat{R}$ values close to 1 and large effective sample sizes [22]. We discuss the results of these models for two cohorts: the 1700-1900 cohort and the GM cohort. Similar results for the other cohorts are included in the supplementary material. All code required to reproduce this analysis is available in the associated Github repository.
Comparing the fitted model for players of different ability.
We first fit the proposed model for standard bullet games as described above with $n=1$, utilising the previous game as the only history. We first examine the key global parameters ($\gamma_{1},\gamma_{2},\mu_{\beta}$), in all four cohorts, with posterior credible intervals for each shown in Figure 3. The main difference is seen in the fixed effects $\gamma_{1}$ and $\gamma_{2}$. The estimated posterior mean of the colour effect ($\gamma_{1}$) is 0.091 and 0.25 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing as white would increase the win probability on average to 0.56 for the GMs but only 0.52 for the 1700-1900s. Similarly, the estimated posterior mean of the rating effect ($\gamma_{2}$) is 0.0037 and 0.0053 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing against an opponent rated 100 points lower increases the win probability on average to 0.63 for GMs and 0.59 for 1700-1900s. This agrees with the common beliefs that such effects have a larger impact in higher rated players [^4] Multiple references are given in [49] and that colour assignment and differences in skill play a larger role in determining the winner of a game at the professional level. In both cohorts the posterior mean value for $\mu_{\beta}$ is centered around 0, indicating that there is little evidence for a global experiential effect. We also include the estimates for each of these parameters for the two intermediate cohorts in Figure 3. We see a clear linear pattern in the estimates of $\gamma_{1}$ and $\gamma_{2}$, corresponding to increasing ability in the cohort. However, for all cohorts, the estimate of $\mu_{\beta}$ clearly centered around 0.
Fig. 3: The 66% and 95% credible intervals of the estimated colour ($\gamma_{1}$), rating ($\gamma_{2}$) and global experiential ($\mu_{\beta}$) effects in the bullet games. The posterior means of the colour and rating effects increase with skill. There is little evidence for a global winner effect in any cohort.
While we see little evidence for a global winner-/loser-effect, we also wish to examine the individual experiential effects and how these vary across focal players. In Figure 4 we take the posterior estimates of $\beta_{j}$ for each player and transform them into the change in predicted win probability. In particular, if all other parameters were fixed, we compute how much the win probability would increase in going from losing the previous game to winning it.
We show histograms and credible intervals of the posterior estimates of the change in win probability for 36 focal players in the 1700-1900 cohort [^5] Four players in the 1700-1900 cohort were excluded due to not playing any 60+0 bullet games. and 25 in the GM cohort. We see that many of these players have changes close to 0. The posterior means for this quantity across these two cohorts range from -0.10 to 0.12. The majority of 95% posterior intervals contain 0 (or being close to 0). These effects are small, and indicate small changes in the win probability based on performance in the previous game. One interesting observation here is that the player with the smallest estimated change in the GM cohort, the username DrNykerstein, is one of the accounts of former world champion Magnus Carlsen. While this result indicates that Carlsen is less likely to win the next game if he had won the previous game than other players, there are several potential factors which could explain this. One possible explanation concerns data quality. Carlsen is widely known for playing in an ``erratic" manner in casual games online, and as seen in live streams, may not be taking such games seriously. In terms of the model used, it is possible that Carlsen is different to other players and (for example) has a winner and loser effect of different sizes. This would not be captured by our current model.
Fig. 4: Interpretation of the estimated player-level experiential effects ($\beta_{j}$ for player $j$). The change in win probability represents the additive boost in win probability a player receives from the experiential effect when changing their previous game from a loss to a win. The 66% and 95% credible intervals are shown. Most players exhibit little evidence for non-zero experiential effects, which results in changes in the win probability that are close to zero. There is a slightly larger proportion of GMs showing evidence of possessing such effects compared to the 1700-1900 cohort.
The Use of Different History
Our proposed model shows little evidence for winner-loser effects above when $n=1$, which corresponds to the result of the previous game alone impacting subsequent performance. However, it is possible that experiential effects could require longer time periods to develop. To investigate this we refit the above model now using $n=10$ for computing the win ratio $\tilde{x}_{ij}^{n}$. It is relatively common for sessions involving more than 10 games to be played consecutively, with more than 20% of all sessions across both cohorts containing at least 11 games. In Figure 5(a) we show the posterior estimates for the three key model parameters for both $n=1$ and $n=10$ for the GM cohort. We note that the posterior distributions for all parameters except for $\tau_{2}$ are essentially identical. In both cases our posterior distribution for $\mu_{\beta}$ is centered around 0, indicating there is no evidence for a global effect. The posterior distribution of $\tau_{2}$ is shifted towards larger values due to the increased variability of $\tilde{x}_{ij}^{n}$ when $n=10$. Individual winner-loser effects (shown in the supplementary material) show similar patterns with increased variability when $n=10$, with a similar proportion of posterior intervals containing 0 in both cases. It does not appear that a longer term history provides additional evidence for the presence of strong or persistent experiential effects.
Comparing Blitz and Bullet Games
The previous analysis has investigated the presence of experiential effects in short bullet games lasting at most two minutes. To consider the possibility that longer competition is required for such effects to form, we repeat the previous analysis now using blitz games. These are games of up to 6 minutes in length. Although these games are longer, it is still common for successive sequences of games. We define a focal player starting another blitz game within 7 minutes of their previous game as games played in the same session. With this definition, two thirds of sessions played in the GM cohort contain more than one game, with almost three quarters of games in the 1700-1900 cohort. As described in Section 2, we have substantially more bullet than blitz games for all players.
We compare the fit of our proposed model with $n=1$ for the GM cohort. The posterior estimates of key parameters are shown in Figure 5(b). The estimated fixed effects $\gamma_{1},\gamma_{2}$ show larger variability in the smaller blitz dataset. In both cases, we see the distribution of the estimated global winner-/loser-effect is centered at 0 with the same variability for both datasets.
(a) Comparing the estimated parameters when $n=1$ vs $n=10$.
(b) Comparing the estimated parameters using bullet or blitz data.
Fig. 5: Comparison of the estimated fixed and global effects in varying settings for the GM cohort. In each case we show 66% and 95% posterior intervals.
In Figure 6 we show the estimated individual winner-loser effects for the members of the GM cohort using both the blitz and bullet data. Some of the distributions appear to shift slightly with the blitz data but all demonstrate larger variability using the blitz data. This is expected, as all play more bullet games (on average 9000 bullet games compared to 1300 blitz games in this data, with an ever bigger difference in other cohorts).
Fig. 6: Winner-loser effects for the GM data, from fitting the proposed model to Bullet and Blitz data. We see considerably larger variability with the smaller blitz dataset.
Overall, we see little evidence for consistent experiential effects from fitting the proposed Bayesian logistic regression model to large volumes of online chess data. Across a variety of formats and time scales for the formation of these effects, we do not see any evidence for such effects across the population of players considered. While there is some evidence for individual players displaying winner-loser effects, these appear to be weak. An important concern with the current analysis, investigating these effects in the context of recent performance, is the potential for high temporal correlation between successive games. In Section 5 we consider this and provide several tools to validate our model.
5 Model Validation
The previous sections demonstrated that, under the proposed hierarchical regression model, there is little evidence for a consistent winner-/loser-effect across the population of online chess players studied. However, our model is very much a simplification of the complex nature of such data. One important consideration which are model does not account for is the temporal nature of this data. Many of the focal players in these datasets have several years of games in the dataset. It is possible that potential experiential effects can change over time, which we would fail to capture in this model. There is the potential for strong temporal dependence in this data which we do not account for in our simplified model. Below we consider several methods to evaluate the fit of our proposed model in the context of these potential issues. We demonstrate that while our model may somewhat suffer from ignoring the clear temporal structure in this data, the current model can capture much of the dependence in this data.
Posterior Predictive Checks
Given the Bayesian model used here, a natural tool to assess our proposed model is to utilise posterior predictive checking. This is a widely used procedure where replicated data under the fitted Bayesian model is compared to the original observed data [22]. Following the notation of [22] we define the posterior predictive distribution as
[…]
where $\theta$ is the vector of parameters in the model. We can then draw replicated data from this distribution, namely we can simulate the result of a given match, given the covariates and each posterior draw of the model parameters. A failure of these replications to match the observed data can identify potential flaws in the model. These posterior predictive checks are often done using some generated quantity from the original dataset. Here, a natural quantity to consider is the number of games won by a focal player in the observed data. An extension of this is to consider the generated quantity corresponding to the rating evolution of a player over replicated datasets. This is the quantity we will use for posterior predictive checking below.
Given the size of our original dataset, creating thousands of posterior predictive replications of each game has large memory requirements. Instead, we will restrict our data to a selection of focal players in the GM cohort. For these players we utilise their 2000 most recent bullet games in the dataset. We will use the first half of this data to estimate the posterior distribution, before then simulating the results of the games in the second half, using each of the draws from the posterior and the value of $\tilde{x}_{ij}^{n}$ based on the simulated result of the previous game with this set of posterior draws. For each set of parameter draws from the posterior we can then compute the estimated ranking under a Glicko-2 rating system, similar to that used by Lichess. We use default values for the constant parameters in this model, which gives similar results to the (unknown) parameters used by Lichess. This gives us posterior predictive draws of the rating of a player over 1000 games, which we can then compare to the true evolution of these scores.
We fit our previous proposed model with $n=1$ on the previous 1000 games to the final 1000 games for the GM cohort. As we use the standard number of chains (4) and sampling iterarions (1000) in CmdStanR, this gives us 4000 posterior draws for each parameter. We then use each set of draws to simulate the results of each game in the last 1000 games. This gives us 4000 replications of the 1000 games played by each focal player and the corresponding evolution of their Glicko-2 score. We show these replications for 2 GMs in Figure 7. While these draws show large uncertainty we see that they capture the overall trend well, in particular capturing the steady increase in rating shown in Figure 7(b). This indicates that our model is able to predict the future temporal evolution of the ability of players well.
(a) Posterior predictive distribution for Magnus Carlsen.
(b) Posterior predictive distribution for Nihal Sarin.
Fig. 7: Posterior predictive distribution of Glicko-2 ratings for 2 GMs. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
Permuting the Game Results
As an alternative approach to assessing the robustness of our model, we consider permuting the data. A similar approach was considered by [11], to examine evidence for a hot hand effect in hockey. For computational considerations, we perform this analysis with a subset of the complete bullet data. We use 12 players from the 1700-1900 cohort and 10 players from the GM cohort, taking the final 1000 games from each. We randomly permute the results within each focal players data 1000 times, giving us 1000 permuted datasets. For each of these we fit our proposed model with $n=1$. We take the posterior mean for each of our parameters from each of these datasets, to construct a permutation distribution for each parameter. In Figure 8(a) we show the distribution of the posterior means under the permuted data, for the 1700-1900 cohort, for the estimated individual winner-/loser-effect for the 20 focal players. The intervals shown correspond to 66% and 95% of these permutation distributions. The estimated posterior means for the true data are shown in red. In each case these permuted datasets give similar estimated effects to those found with the real data, and all estimated effects agree with these null distributions.
We show the same analysis for the 10 players in the GM cohort in Figure 8(b). We see similar distributions for the individual effects under these permutations, although we see that the permutation intervals do not describe the true effects as well as in the 1700-1900 cohort. This is a general pattern we see where it appears that our model is less appropriate as the ability of players increases, and we see this as the rating increases. The estimated effects for the true data are similar to those estimated under the permutations. We note that when we repeat this analysis for the fixed effects $\gamma_{1}$ and $\gamma_{2}$ the distribution of these effects under the permutations is clearly different from the estimated effects, for all cohorts, as would be expected.
(a) Permutation estimates for 20 players in 1700-1900 Cohort.
(b) Permutation estimates for 10 players in the GM cohort.
Fig. 8: Posterior mean estimates for winner/loser-effects after permuting the game data. For each the true posterior mean estimate is shown in red, with the intervals representing 66% and 95% of the permutation distributions.
6 Summary, Future Directions
In this work we have proposed a Bayesian hierarchical logistic regression model to investigate the presence of experiential effects in online chess. having fit this model to cohorts of players of varying ability we find that there is no evidence for the presence of global winner/loser-effects, as commonly defined in the literature. However, a small proportion of individual players from all cohorts show evidence of possessing such effects. This is consistent across multiple formats of chess and the number of previous games that this effect could persist over. While we do not believe that our model perfectly describes the dynamics of chess, our model checking procedures indicate that our model appears reasonable and captures many of the key characteristics of this data. Our results indicate that a more flexible model may be required for players of higher ability, such as GMs. This would need to be investigated further, and could be related to the larger impact of the covariates (color and rating difference) in games between better players.
There are many potential future directions which could expand on this work. In shorter chess games it is common to attempt to ``flag" your opponent, aiming to beat them by forcing them to run out of time. A large proportion of bullet games are won this way (24% of bullet games in the GM cohort) and results of this form may have a different impact on future performance than other endings. We have also not attempted to separate out winner and loser effects or allow them to have different magnitudes in the current formulation. The current model also does not consider draws (which are rare in this data) and the potential impact they could have on future performance.
\@noitemerr
Please insert acknowledgments of the assistance of colleagues or similar notes of appreciation here.
\@noitemerr
Please insert information concerning research grant support here (institution and grant number). Please provide for each funder the funder’s DOI according to https://doi.crossref.org/funderNames?mode=list.
Appendix: Additional Analyses
Here we include some additional model fitting results which were omitted from the main text. In particular, we show selected model estimates and model diagnostics for the other cohorts considered. We also investigate the effect of using varying numbers of games for the historic influence.
Examining Winner-/Loser-Effects across Cohorts
In Figure A1 we show the posterior distributions of individual $\beta$ parameters for each player in the 2000-2200 and 2300-2500 cohorts. In each case we see similar results to those shown in the main text. The majority of posterior distributions for $\beta_{j}$ contain zero, with a small number of players showing positive or negative estimates. This behaviour is common across all ranges of rating we considered.
(a) Players in the 2000-2200 cohort.
(b) Players in the 2300-2500 cohort.
Fig. A1: The estimated change in win probability for the other two cohorts we consider.
Posterior Predictive Checking for Other Cohorts
In Figure A2 we demonstrate further posterior predictive checks using players from cohorts not included in the main text. We again construct the posterior predictive distributions of the Glicko-2 rating evolution. We show sample players in the 1700-1900, 2000-2200 and 2300-2500 cohorts. In each case we see that the predictive distribution appears to match the general pattern over future games, indicating our model is able to predict the temporal nature of the ranking evolution.
(a) Posterior predictive distribution for "APlayerfromEarth" (1700-1900 Cohort).
(b) Posterior predictive distribution for "Rerereggie" (2000-2200 Cohort).
(c) Posterior predictive distribution for "Lobovanvliet" (2300-2500 Cohort).
Fig. A2: Posterior predictive distribution of Glicko-2 ratings for 3 players in different rating cohorts. We show the true Lichess rating in red, with 4000 draws of the predicted rating evolution. The mean of these draws is shown in the solid black line.
Investigating the Role of $n$
In the main text we fit our model with $n=1$ and $n=10$, considering the previous performance over the previous game and the previous 10 games respectively. Given that many playing sessions contain less than 10 games, a natural question is whether $n=10$ is too long a time scale for potential experiential effects to persist. In Figure A3 we show the key global parameters $\gamma_{1},\gamma_{2},\mu_{\beta}$ for the 1700-1900 and GM cohorts, for $n=1,5,10$. In each case we see that the estimates of these parameters are essentially identical, which is also the case for the individual parameters $\beta_{j}$.
(a) 1700-1900 Cohort
(b) GM Cohort
Fig. A3: Posterior estimates for the key global parameters $\gamma_{1},\gamma_{2},\mu_{\beta}$ as we vary $n$ in our fitted model.
Player Effects
Finally, we show some of the estimated player effects which we have excluded from the main text. In Figure A4 we show the posterior estimates for each $\alpha_{j}$ for the 1700-1900 and GM cohort. For the lower ability players many of these effects appear to be negative. This agrees with intuition that these players are likely to win less than half of their games. For the GM cohort we see some strong positive player effects, including for Magnus Carlsen.
(a) Player Effects for the 1700-1900 cohort.
(b) Player Effects for the GM cohort.
Fig. A4: Estimated Player Effects for the 1700-1900 and GM cohorts.